HanAkoSolo (jamierox4ev3r):

What are the foci of the ellipse given by the equation: 16x^2 + 9y^2 + 64x + 108y + 244 = 0? A) (–2, –6 ±square root of 7) B) (3, –6) and (–7, –6) C) (–2, –1) and (–2, –11) D) (-2 ±square root of 7 , –6) How would I get the equation to the regular formula of ellipses? like x^2/a^2 + y^2/b^2 = 1. A question from @Loujoelou , since his internet is currently slow.

HanAkoSolo (jamierox4ev3r):

@litchlani can u help with this one?

OpenStudy (anonymous):

you make it look like \[\frac{(x-h)^2}{a^2}+\frac{(y-k)^2}{b^2}=1\] by completing the square

OpenStudy (anonymous):

it won't look the way you wrote because the center is not \((0,0)\)

OpenStudy (anonymous):

first group it up like this : \[16(x^{2}+4x)+16(y^{2}+y)+244=0\] then complete the square for the thing in the brackets

OpenStudy (anonymous):

that wrong wait

HanAkoSolo (jamierox4ev3r):

? if its wrong please erase the comment, don't want to get confused between right and wrong

OpenStudy (anonymous):

second one should be \[9(y^2+12y)\] sry

HanAkoSolo (jamierox4ev3r):

oh okay..

OpenStudy (anonymous):

now u can complete the square for the things in the brackets , you know how to do it ?

HanAkoSolo (jamierox4ev3r):

yeah should be good thank you :)

HanAkoSolo (jamierox4ev3r):

thanks to both of you i should be able to find the foci by myself